OpenStudy (anonymous):

What are the roots of the following polynomial equation? (x + 5)(x + 5i)(x - 5i) = 0

OpenStudy (amoodarya):

(x-x1)(x-x2)(x-x3)...(x-xk)=0 roots are x1, x2,x3,... because x-x1=0 ---> x=x1 and so on so (x + 5)(x + 5i)(x - 5i) = 0 --> x+5=0 x=-5 x+5i=0 x=-5i x-5i=0 x=5i but 5 is a real root 5i ,-5i are imaginary root (complex roots)

OpenStudy (anonymous):

ok so are my roots -5, 5i and -5i???

OpenStudy (anonymous):

is anyone there?

OpenStudy (amoodarya):

yes -5, 5i and -5i